OpenStudy (anonymous):

Troy and his wife Angela put $5,000.00 of their savings into a bank account earning 7.25% interest. How much will Trey and Angela have in all after 3 years?

OpenStudy (anonymous):

metal is going to be given show your work tho possible answers are $90.63 $362.50 $1,087.50 $6,087.50

OpenStudy (kash_thesmartguy):

Interest=5000*0.0725*3 Interest=$1087.5 Your Answer= 1087.5+5000=$6087.5

OpenStudy (anonymous):

thanks @Kash_TheSmartGuy